OpenStudy (anonymous):

what are the zeros of the related function to this equation x^2-4x-21=0

OpenStudy (welshfella):

To factor this you need 2 numbers, whose product is -21 ans whose sum is -4

OpenStudy (cwrw238):

it factors to (x + a)(x + b) = 0 where a and b are the 2 numbers

OpenStudy (anonymous):

so the zeros are -21 and 4 or just 2?

OpenStudy (cwrw238):

no a and b have to be found first ab = -21 a + b = -4 can you solve this system of equations?

OpenStudy (cwrw238):

a and b are the 2 numbers welshfella talked about

OpenStudy (anonymous):

a+b+-4 -7+3?

OpenStudy (cwrw238):

right -7*3 = -21 and -7 + 3 = -4

OpenStudy (cwrw238):

so we have (x - 7)(x + 3) = 0 now it easy to find the zeroes

OpenStudy (cwrw238):

can you solve that?

OpenStudy (cwrw238):

you have 2 expressions multiplied together to give zero therefore either one of them must equal zero so we have x - 7 = 0 and x + 3 = 0 solve these 2 equations and you have your answer

OpenStudy (anonymous):

x=7 & x=3?

OpenStudy (cwrw238):

not quite

OpenStudy (cwrw238):

one of them should be negative

OpenStudy (anonymous):

-7?

OpenStudy (cwrw238):

x + 3 = 0 add -3 to both sides

OpenStudy (anonymous):

-3... my bad

OpenStudy (cwrw238):

x + 3 - 3 = -3 x = ?

OpenStudy (cwrw238):

yes answer is {-3,7}

OpenStudy (cwrw238):

- this is called the solution set

OpenStudy (anonymous):

im really lost im sorry

OpenStudy (cwrw238):

which part are you lost on?

OpenStudy (anonymous):

the question is what are the ZEROS of the related function im just confused at finding the zeros I guess...

OpenStudy (cwrw238):

the zeros are the values of x which make the function equal to zero That is why you write x^2 - 4x - 21 = 0. They correspond to the points where the function cuts the x-axis ( where the function = 0)

OpenStudy (cwrw238):

the graph looks a bit like this